Evaluate the following integrals:

  1. ∫ 𝑑 π‘π‘œπ‘ 2𝑑 𝑑𝑑
  2. ∫ π‘₯^3 𝑙𝑛π‘₯ 𝑑π‘₯
  3. ∫ π‘₯^3 𝑒^3π‘₯ 𝑑π‘₯
  4. ∫ 𝑠𝑖𝑛(𝑙𝑛π‘₯) 𝑑π‘₯
  5. βˆ«π‘ π‘–π‘›^βˆ’1 π‘₯ 𝑑π‘₯
  6. ∫ 𝑦^3 𝑒^𝑦^2 𝑑𝑦
1
Expert's answer
2021-05-14T14:58:37-0400

1.

"\int t cos (2t) dt"

Integrate by parts


"{\int}\mathtt{f}\mathtt{g}' = \mathtt{f}\mathtt{g} - {\int}\mathtt{f}'\mathtt{g}\\
f=t, g'=\cos\left(2t\right)\\
f'=1\, , g=\dfrac{\sin\left(2t\right)}{2}\\"

Now solving:


"{\displaystyle\int}\dfrac{\sin\left(2t\right)}{2}\,\mathrm{d}t\\
\text{Substitute } u=2t \rightarrow \frac{du}{dt} =2 \rightarrow dt= \dfrac{1}{2}du:\\
= \frac{1}{4}\int \sin(u)du\\
=-cos(u)"


Plugging in solved integrals:


"\frac{1}{4}\int \sin(u)du\\
=-\frac{cos u}{4}\\
\text{ Recall that } u=2t\\
\therefore -\frac{cos (2t)}{4}\\"

Plug in solved integrals,


"\dfrac{t\sin\left(2t\right)}{2}-{\displaystyle\int}\dfrac{\sin\left(2t\right)}{2}\,\mathrm{d}t\\
=\dfrac{t\sin\left(2t\right)}{2}+\dfrac{\cos\left(2t\right)}{4}"

Thus:


"{\displaystyle\int}t\cos\left(2t\right)\,\mathrm{d}t =\dfrac{t\sin\left(2t\right)}{2}+\dfrac{\cos\left(2t\right)}{4}+C\\
=t\cos\left(t\right)\sin\left(t\right)+\dfrac{\cos^2\left(t\right)}{2}+C"

2.

"\int x^3 \ln x\; dx"

"\text{Integrate by parts:} \\\,\\
\int \mathrm{fg}^{\prime}=\mathrm{fg}-\int \mathrm{f}^{\prime}{\mathrm{g}}\\
f=\ln (x), g^{\prime}=x^{3}\\
\quad f^{\prime}=\frac{\downarrow_{\text {steps }}}{x}, \quad g=\frac{\downarrow_{\text {steps }}}{4}\\
=\frac{x^{4} \ln (x)}{4}-\int \frac{x^{3}}{4} d x\\
\text{Now solving:}\\\,\\
\int \frac{x^{3}}{4} d x\\
\text{Apply linearity:}\\\,\\
=\frac{1}{4} \int x^{3} d x\\
\text{Now solving:}\\
\int x^{3} d x\\
\text{Apply power rule:}\\
\begin{aligned} \int x^{n} d x=& \frac{x^{n+1}}{n+1} \text { with } n=3: \\ &=\frac{x^{4}}{4} \end{aligned}\\\,\\
\text{Plug in solved integrals:}\\
\frac{1}{4} \int x^{3} d x\\
=\frac{x^{4}}{16}\\\,\\
\text{Plug in solved integrals:}\\
\frac{x^{4} \ln (x)}{4}-\int \frac{x^{3}}{4} d x\\
=\frac{x^{4} \ln (x)}{4}-\frac{x^{4}}{16}\\\,\\
\text{Thus}\\
\int x^{3} \ln (x) d x=\frac{x^{4} \ln (x)}{4}-\frac{x^{4}}{16}+C\\
=\frac{x^{4}(4 \ln (x)-1)}{16}+C"


3.


"{\displaystyle\int}x^3\mathrm{e}^{3x}\,\mathrm{d}x"

Using integration by parts


"\int x^{3} \mathrm{e}^{3 x} \mathrm{d} x\\\,\\
\text{Integrate by parts:} \\
\int \mathrm{fg}^{\prime}=\mathrm{fg}-\int \mathrm{f}^{\prime} \mathrm{g}\\
\mathrm{f}=x^{3}, \mathrm{g}^{\prime}=\mathrm{e}^{3 x} \\
\mathrm{f}^{\prime}=3 x^{2}, \mathrm{~g}=\frac{\mathrm{e}^{3 x}}{3} \\
=\frac{x^{3} \mathrm{e}^{3 x}}{3}-\int x^{2} \mathrm{e}^{3 x} \mathrm{d} x\\
\,\\
\text{Now solving:}\\
\int x^{2} \mathrm{e}^{3 x} \mathrm{d} x\\\,\\
Integrate by parts: \\
\int \mathrm{fg}^{\prime}=\mathrm{fg}-\int \mathrm{f}^{\prime} \mathrm{g}\\
\mathrm{f}=x^{2}, \mathrm{~g}^{\prime}=\mathrm{e}^{3 x} \\
\mathrm{f}^{\prime}=2 x, \mathrm{~g}=\frac{\mathrm{e}^{3 x}}{3} \\
=\frac{x^{2} \mathrm{e}^{3 x}}{3}-\int \frac{2 x \mathrm{e}^{3 x}}{3} \mathrm{d} x\\
\,\\
\text{Now solving:}
\\
\int \frac{2 x \mathrm{e}^{3 x}}{3} \mathrm{d} x
\\\,\\
\text{Apply linearity:}

=\frac{2}{3} \int x \mathrm{e}^{3 x} \mathrm{~d} x\\

\text{Now solving:}\\
\int x \mathrm{e}^{3 x} \mathrm{~d} x
\\\,\\
\text{Integrate by parts:} \\
\int \mathrm{fg}^{\prime}=\mathrm{fg}-\int \mathrm{f}^{\prime} \mathrm{g}\\
\mathrm{f}=x, \mathrm{g}^{\prime}=\mathrm{e}^{3 x} \\
\mathrm{f}^{\prime}=1, \mathrm{g}=\frac{\mathrm{e}^{3 x}}{3} \\
=\frac{x \mathrm{e}^{3 x}}{3}-\int \frac{\mathrm{e}^{3 x}}{3} \mathrm{d} x
\\\,\\
\text{Now solving:}
\\
\int \frac{\mathrm{e}^{3 x}}{3} \mathrm{d} x \\
\text { Substitute } u=3 x \longrightarrow \frac{\mathrm{d} u}{\mathrm{d} x}=3 \longrightarrow \mathrm{d} x=\frac{1}{3} \mathrm{d} u: \\
=\frac{1}{9} \int \mathrm{e}^{u} \mathrm{d} u
\\\,\\
\text{Now solving:}
\\
\int \mathrm{e}^{u} \mathrm{d} u
\\\,\\
\text{Apply exponential rule:}\\
\int \mathrm{a}^{u} \mathrm{~d} u= \frac{\mathrm{a}^{u}}{\ln (\mathrm{a})} \text { with } \mathrm{a}=\mathrm{e}: \\
=\mathrm{e}^{u}\\\,\\
\text{Plug in solved integrals}:
\\
\frac{1}{9} \int \mathrm{e}^{u} \mathrm{~d} u \\
=\frac{\mathrm{e}^{u}}{9}\\
\\
\text{Undo substitution} u=3 x :
\\
=\frac{\mathrm{e}^{3 x}}{9}
\\
\text{Plug in solved integrals}:
\\
\frac{x \mathrm{e}^{3 x}}{3}-\int \frac{\mathrm{e}^{3 x}}{3} \mathrm{~d} x \\
=\frac{x \mathrm{e}^{3 x}}{3}-\frac{\mathrm{e}^{3 x}}{9}
\\
\text{Plug in solved integrals}:
\\
\frac{2}{3} \int x \mathrm{e}^{3 x} \mathrm{~d} x \\
=\frac{2 x \mathrm{e}^{3 x}}{9}-\frac{2 \mathrm{e}^{3 x}}{27}
\\
\text{Plug in solved integrals}:\\
\frac{x^{2} \mathrm{e}^{3 x}}{3}-\int \frac{2 x \mathrm{e}^{3 x}}{3} \mathrm{~d} x\\
=\frac{x^{2} \mathrm{e}^{3 x}}{3}-\frac{2 x \mathrm{e}^{3 x}}{9}+\frac{2 \mathrm{e}^{3 x}}{27}\\
\text{Plug in solved integrals}:\\
\frac{x^{3} \mathrm{e}^{3 x}}{3}-\int x^{2} \mathrm{e}^{3 x} \mathrm{~d} x\\
=\frac{x^{3} \mathrm{e}^{3 x}}{3}-\frac{x^{2} \mathrm{e}^{3 x}}{3}+\frac{2 x \mathrm{e}^{3 x}}{9}-\frac{2 \mathrm{e}^{3 x}}{27}\\
\text{Thus:}\\
\int x^{3} \mathrm{e}^{3 x} \mathrm{~d} x=\\\frac{x^{3} \mathrm{e}^{3 x}}{3}-\frac{x^{2} \mathrm{e}^{3 x}}{3}+\frac{2 x \mathrm{e}^{3 x}}{9}-\frac{2 \mathrm{e}^{3 x}}{27}+C\\
=\frac{\left(9 x^{3}-9 x^{2}+6 x-2\right) \mathrm{e}^{3 x}}{27}+C"

4.


"\int{\sin{\left(\ln{\left(x \right)} \right)} d x}"

"\begin{array}{c}
\text { Substitute } u=\ln (x) \longrightarrow \frac{\mathrm{d} u}{\mathrm{~d} x}=\frac{1}{x}=\mathrm{e}^{u} \longrightarrow \mathrm{d} x=x \mathrm{~d} u, \text { use: } \\
=\int \mathrm{e}^{u} \sin (u) \mathrm{d} u
\end{array}
\\
\text{ We will integrate by parts twice in a row: } \\
\int \mathrm{fg}^{\prime}=\mathrm{fg}-\int \mathrm{f}^{\prime} \mathrm{g}\\

\text{ First time: }
\\ \mathbf{f}=\sin (u), \quad \mathrm{g}^{\prime}=\mathrm{e}^{u} \\
\begin{array}{c}\\
\mathbf{f}^{\prime}=\cos (u), \mathrm{g}=\mathrm{e}^{u} \\
=\mathrm{e}^{u} \sin (u)-\int \mathrm{e}^{u} \cos (u) \mathrm{d} u
\end{array}
\\

\text{ Second time:} 
\\
\begin{array}{c}\\
\mathrm{f}=\cos (u), \quad \mathrm{g}^{\prime}=\mathrm{e}^{u} \\
\mathrm{f}^{\prime}=-\sin (u), \mathrm{g}=\mathrm{e}^{u}: \\
=\mathrm{e}^{u} \sin (u)-\left(\mathrm{e}^{u} \cos (u)-\int-\mathrm{e}^{u} \sin (u) \mathrm{d} u\right)
\end{array}\\

\text { Apply linearity: }\\
=\mathrm{e}^{u} \sin (u)-\left(\mathrm{e}^{u} \cos (u)+\int \mathrm{e}^{u} \sin (u) \mathrm{d} u\right)\\
\text{ The integral } \int \mathrm{e}^{u} \sin (u) \mathrm{d} u\\
 \text{ appears again on the right side of the equation, we can solve for it: }\\

=\frac{\mathrm{e}^{u} \sin (u)-\mathrm{e}^{u} \cos (u)}{2}\\
\text{ Undo substitution } u=\ln (x), \text { use: }
\\
\begin{array}{c}\mathrm{e}^{\ln (x)}=x \\ =\frac{x \sin (\ln (x))-x \cos (\ln (x))}{2}\end{array}
\\
\text{Thus }\\
\int \sin (\ln (x)) \mathrm{d} x
=\frac{x \sin (\ln (x))-x \cos (\ln (x))}{2}+C\\
\int \sin (\ln (x)) \mathrm{d} x=\frac{x(\sin (\ln (x))-\cos (\ln (x)))}{2}+C\\"

5.


"\int \arcsin (x) \mathrm{d} x\\
\text{ Integrate by parts: }\\
\int \mathrm{fg}^{\prime}=\mathrm{fg}-\int \mathrm{f}^{\prime} \mathrm{g}\\

f=\arcsin (x), g^{\prime}=1 \\
{f^{\prime}}=\frac{1}{\sqrt{1-x^{2}}}, g=x \\
=x \arcsin (x)-\int \frac{x}{\sqrt{1-x^{2}}} \mathrm{~d} x\\

\text{ Now solving: }\\
\int \frac{x}{\sqrt{1-x^{2}}} d x\\
\text { Substitute } u=1-x^{2} \longrightarrow \frac{\mathrm{d} u}{\mathrm{~d} x}=-2 x \longrightarrow \mathrm{d} x=-\frac{1}{2 x} \mathrm{~d} u \text { : } \\ =-\frac{1}{2} \int \frac{1}{\sqrt{u}} \mathrm{~d} u\\

\text{ Now solving: }\\
\int \frac{1}{\sqrt{u}} \mathrm{~d} u

\text{ Apply power rule: }\\

\int u^{\mathrm{n}} \mathrm{d} u=\frac{u^{\mathrm{n}+1}}{\mathrm{n}+1} \text { with } \mathrm{n}=-\frac{1}{2}: \\
=2 \sqrt{u}
\\

\text{ Plug in solved integrals: }\\
-\frac{1}{2} \int \frac{1}{\sqrt{u}} \mathrm{~d} u\\
=-\sqrt{u}
\\
\text{ Undo substitution } \boldsymbol{u}=\mathbf{1}-\boldsymbol{x}^{2}\\
=-\sqrt{1-x^{2}}\\
\text{ Plug in solved integrals :}\\
x \arcsin (x)-\int \frac{x}{\sqrt{1-x^{2}}} \mathrm{~d} x\\
=x \arcsin (x)+\sqrt{1-x^{2}}\\
\text{ Thus: }\\
\int \arcsin (x) \mathrm{d} x
=x \arcsin (x)+\sqrt{1-x^{2}}+C"

6.


"\int y^{3} \mathrm{e}^{y^{2}} \mathrm{~d} y \\
\text { Substitute } u=y^{2} \longrightarrow \frac{\mathrm{d} u}{\mathrm{~d} y}=2 y \longrightarrow \mathrm{d} y=\frac{1}{2 y} \mathrm{~d} u \text { : } \\= \frac{1}{2} \int u \mathrm{e}^{u} \mathrm{~d} u\\

\text{ Now solving: }
\int u \mathrm{e}^{u} \mathrm{~d} u\\

\text{ Integrate by parts: } \\
\int \mathrm{fg}^{\prime}=\mathrm{fg}-\int \mathrm{f}^{\prime} \mathrm{g}\\
\begin{array}{l}
\mathrm{f}=u, \quad \mathrm{~g}^{\prime}=\mathrm{e}^{u} \\
\mathrm{f}^{\prime}=1, \mathrm{~g}=\mathrm{e}^{u} \\
=u \mathrm{e}^{u}-\int \mathrm{e}^{u} \mathrm{~d} u
\end{array}\\
\text{ Now solving: } \int \mathrm{e}^{u} \mathrm{~d} u\\

\text{ Apply exponential rule: }\\
 \int a^{u} \mathrm{~d} u= \frac{\mathrm{a}^{u}}{\ln (\mathrm{a})} \text { with } \mathrm{a}=\mathrm{e}: \\ =\mathrm{e}^{u} \\

\text{ Plug in solved integrals: }\\
u \mathrm{e}^{u}-\int \mathrm{e}^{u} \mathrm{~d} u \\
=u \mathrm{e}^{u}-\mathrm{e}^{u}\\

\text{ Plug in solved integrals: }\\
\frac{1}{2} \int u \mathrm{e}^{u} \mathrm{~d} u \\
= \frac{u \mathrm{e}^{u}}{2}-\frac{\mathrm{e}^{u}}{2}\\

\text{ Undo substitution }\boldsymbol{u}=\boldsymbol{y}^{2} :\\
=\frac{y^{2} \mathrm{e}^{y^{2}}}{2}-\frac{\mathrm{e}^{y^{2}}}{2}\\

\text{ Thus: }\\
\int y^{3} \mathrm{e}^{y^{2}} \mathrm{~d} y =\frac{y^{2} \mathrm{e}^{y^{2}}}{2}-\frac{\mathrm{e}^{y^{2}}}{2}+C\\
\int y^{3} \mathrm{e}^{y^{2}} \mathrm{~d} y=\frac{\left(y^{2}-1\right) \mathrm{e}^{y^{2}}}{2}+C"